if line L bisects segment AB at point m, and if AM=4x and MB=5x-4, find AB
Answer by ikleyn(52835)

.
 
    4x = 5x - 4

    0  =  x - 4

    x = 4

    AB = AM + MB = = 2 * AM = 2 * (4*4) = 2 * 16 = 32.    ANSWER

Solved.